dootdoot:

How do you find the roots of the following equation? I have my exam and i wanna know the way to solve this myself or with desmos.

dootdoot:

Aj3aiden:

Can you factor it??????

Echofire:

Try Symbolab, it has free steps and graphs. You can even do chemistry and math up to like calculus 4 on there.

imqwerty:



dootdoot:

how would i solve this guy

Echofire:

Alright, so what I mean by "Symbolab" was symbolab.com.

dootdoot:

@echofire wrote:
Alright, so what I mean by "Symbolab" was symbolab.com.
ah, thank you!

dootdoot:

I think i got it for this one

@dootdoot wrote:
how would i solve this guy
I just need to plug in numbers for both guys until they equal the same thing

surjithayer:

\[9^{x^2-4x}=3^{-6}\] \[(3^2)^{x^2-4x}=3^{-6}\] \[3^{2(x^2-4x)}=3^{-6}\] \[2(x^2-4x)=-6\] \[x^2-4x=-3\] \[x^2-4x+3=0\] \[x^{2}-x-3x+3=0\] x(x-1)-3(x-1)=0 (x-1)(x-3)=0 x=1,3